Cost Accounting

The following manufacturing costs were incurred by the Miracle Mile Company in 2019:
 





Direct materials

$

225,000


Direct labor


350,000


Manufacturing overhead


470,000


 
 


These costs were incurred to produce 50,000 units of product. Variable manufacturing overhead was 70% of the direct materials cost.


In 2020, the direct material and variable overhead costs per unit will increase by 12%, but the direct labor costs per unit are not expected to change. Fixed manufacturing costs are expected to increase by 8%.

Required:

(a.) Prepare a cost estimate for an activity level of 40,000 units of product in 2020.

(b.) Determine the total product costs per unit for 2019 and 2020.  
Answer #1

(a)

For 2019

Variable manufacturing overheads = 70% of direct material cost

= 225,000 x 70%

= $157,500

Fixed manufacturing overheads = Total manufacturing overheads - Variable manufacturing overheads

= 470,000 - 157,500

= $312,500

Direct material cost per unit = Direct material cost/Number of units

= 225,000/50,000

= $4.5

Direct labor cost per unit = Direct labor cost/Number of units

= 350,000/50,000

= $7

Variable manufacturing overhead per unit = Variable manufacturing overheads/Number of units

= 157,500/50,000

= $3.15

Fixed manufacturing overhead per unit = Fixed manufacturing overheads/Number of units

= 312,500/50,000

= $6.25

For 2020

In 2020, Direct material cost per unit and Variable manufacturing overhead per unit will increase by 12%

Direct material cost per unit = 4.5 x 112%

= $5.04

Variable manufacturing overhead per unit = 3.15 x 112%

= $3.528

Fixed manufacturing overheads are expected to increase by 8%

Fixed manufacturing overheads = 312,500 x 108%

= $337,500

Cost estimate for 2020

Direct material cost (40,000 x 5.04)201,600
Direct labor cost (40,000 x 7)280,000
Variable manufacturing overhead (40,000 x 3.528)141,120
Fixed manufacturing overhead337,500
Total product cost$960,220

b)

Total product cost per unit for 2019 = Direct material cost per unit + Direct labor cost per unit + Variable manufacturing overhead per unit + Fixed manufacturing overhead per unit

= 4.5 + 7 + 3.15 + 6.25

= $20.9

Total product cost per unit for 2020 = Total product cost/Number of units

= 960,220/40,000

= $24


answered by: ANURANJAN SARSAM